OpenStudy (anonymous):

Write the quadratic equation in factored form. Be sure to write the entire equation. x2 - 5x - 24 = 0

jimthompson5910 (jim_thompson5910):

Hint: Find two numbers that multiply to -24 (the last number) and add to -5 (the middle coefficient)

OpenStudy (anonymous):

do i need to find the GCF of -24

jimthompson5910 (jim_thompson5910):

no, list all the ways to multiply to -24 1*(-24) = -24 2*(-12) = -24 3*(-8) = -24 4*(-6) = -24 (-1)*(24) = -24 (-2)*(12) = -24 (-3)*(8) = -24 (-4)*(6) = -24 Which of these numbers also add to -5 ?

OpenStudy (anonymous):

so I just pick which one i want to use and add -5 to it

jimthompson5910 (jim_thompson5910):

no, of that list, which two numbers add to -5 ? Ex: 1 and -24 multiply to -24, but they do NOT add to -5

OpenStudy (anonymous):

so I got 3*(-8) = -24

jimthompson5910 (jim_thompson5910):

good, 3 + (-8) = -5

jimthompson5910 (jim_thompson5910):

So x^2 - 5x - 24 = 0 factors to (x+3)(x-8) = 0

OpenStudy (anonymous):

Thank you

jimthompson5910 (jim_thompson5910):

np